# SpecServe: Efficient and SLO-Aware Large Language Model Serving with Adaptive Speculative Decoding

### v1

@@ 7.3 Ablation Study
As shown in Table 2, we conducted an ablation study of the performance of different dynamic methods on real-world production traces, exploring the sources of speedup for each method across three dimensions: average speculative length, average accepted draft tokens, and average draft token acceptance rate. Specifically, SpecServe- solely utilizes an adaptive drafter, without employing the confidence prior verifier.
Benefits of adaptive drafter. Effective control of speculative length has achieved an additional acceleration ratio of 6% to 55%. We observe that SpecServe- can achieve an average draft token acceptance rate that is twice as high as that based on the threshold-based method, and the higher the acceptance rate, the greater the additional speedup achieved by SpecServe-. This indicates that the speedup of SpecServe- primarily benefits from its ability to efficiently convert draft tokens into accepted tokens. Crucially, the adaptive drafter enhances the average draft token acceptance rate by adopting longer speculative lengths for batches that are likely to accept more tokens. Unlike the threshold-based method that only control speculative length by comparing confidence scores, the adaptive drafter employs a model based on confidence scores to estimate the expected number of accepted tokens per batch, achieving more precise control of speculative length.
Benefits of confidence prior verifier. Request-level control of speculative length further enhances the acceleration ratio by an additional 2% to 5%. SpecServe possesses the highest average draft token acceptance rate among all dynamic methods, which is precisely why it achieves the highest speedup. Compared to SpecServe-, SpecServe slightly affects the average number of accepted draft tokens (a reduction of only 10% of the average speculative length) while reducing the average speculative length, resulting in an increase in the average draft token acceptance rate by 2% to 4%. This indicates that for every ten draft tokens eliminated by the confidence prior verifier, only one could have been accepted, demonstrating the effectiveness of the draft token elimination strategy.
7.4 Behavior Analysis
@@
Figure 12: Fluctuations in batch sizes and speculative lengths under different methods and traces.
@@
As illustrated in Fig. 12, we analyzed the behavior of optimal vanilla speculative decoding (with a speculative length of 3), threshold-based dynamic speculative decoding, and SpecServe across traces from three real production environments. Within a 10-minute observation period, we monitored the current batch size and speculative length every 5 seconds.
Effective speculative length control. Across different production traces, SpecServe’s speculative length generally exhibits an inverse relationship with batch size changes, i.e., as batch size increases, speculative length decreases. In contrast, the speculative length of threshold-based dynamic speculative decoding is unrelated to batch size, often utilizing the maximum speculative length of 8, resulting in a higher average speculative length, as observed in §7.3. On the other hand, the preset length of vanilla speculative decoding cannot adjust speculative length, and its experimental outcomes rely entirely on empirical data. Therefore, through effective and fine-grained speculative length control, SpecServe can adapt to varying request patterns and system configurations, making it highly suitable for integration into actual production environment service systems.
System behaviors. SpecServe generally exhibits lower batch size peaks in behavior analysis, indicating superior performance in handling system request backlogs. Batch size essentially reflects the system’s request backlog volume. Ideally, the system aims to maintain a constant request backlog volume, meaning the number of requests processed by the system within a specific timeframe equals the number of requests received, without resource wastage or shortage. However, in real-world scenarios, especially during sudden high request rates (e.g., the high request rate burst starting at the 410-th second in trace 1), the request backlog volume may sharply increase, placing pressure on system operation. In such instances, SpecServe’s batch size peak is only 43% of that achieved by the threshold-based method and 88% of the optimal vanilla speculative decoding. This effectiveness is primarily attributed to SpecServe’s effective speculative length control mechanism. Faced with a large volume of requests, SpecServe can automatically switch to autoregressive decoding mode, maintaining lower end-to-end latency and effectively avoiding request backlogs. Overall, SpecServe’s stable latency reduction feature achieves smoother system operation, effectively addressing sudden request spikes in real scenarios.
8 Related Work
Speculative decoding solutions. The efficiency of speculative decoding is influenced not only by the speculative length but also depends on the execution efficiency of the draft model and the probability of the draft tokens being accepted by the target model. Hence, developing draft models that are both efficient and similar to the target model has become a focus of speculative decoding research. Previous studies have mainly adopted two approaches to designing draft models: (1) Independent draft models, which involve introducing a high-efficiency independent draft model to generate draft tokens. This model could be a variant from the same series as the target model [18, 33], a fine-tuned model [26, 15, 23, 45], or based on databases from other textual resources [13]; (2) Self-draft models, which adjust the structure of the target model to enable it to rapidly generate multiple draft tokens [4, 19, 11, 42, 20]. Moreover, considerable work has applied speculative decoding to specific scenarios, such as tasks involving small batches of ultra-long sequences [32] and large batches of medium-length sequences [6], designing specific efficient draft models. The diversity of draft models has led to the flourishing development of the speculative decoding, and SpecServe can adaptively control the speculative length for any type of draft model with fine granularity.
AdaServe [21] and our work are contemporaneous, with the following differences: (1)AdaServe focuses on providing SLO-customized services for requests, while SpecServe is committed to achieving stable acceleration guarantees for LLM inference under dynamic scenarios and strict SLO constraints. (2)AdaServe only requires the expected value of TPOT to meet the SLO, whereas SpecServe proposes that the worst-case scenario of TPOT (i.e., the upper limit of the speculative decoding step time) must always strictly comply with the SLO constraints. (3)AdaServe primarily targets tree-based speculative decoding, while SpecServe specifically optimizes for sequential speculative decoding.
Resource-efficient LLM serving. In addition to speculative decoding, existing LLM serving systems propose other solutions to improve resource efficiency. For instance, Orca [38] leverages continuous batching to reduce request queuing delays, thereby achieving efficient GPU resource utilization; vLLM [17] constructs a memory-efficient service system based on PagedAttention; while Llumnix [31] focuses on multi-tenant scenarios, enabling efficient cross-node resource utilization through runtime rescheduling across multiple model instances. Furthermore, the disaggregated architecture [29, 30, 44] has also become a focal point, where the disaggregation of prefill and decoding The separation of prefill and decoding can optimize resource utilization and achieve a balance between TTFT and TPOT. SpecServe enables speculative decoding to serve as an efficient and stable decoding solution, forming an orthogonal synergy with the resource-efficient techniques of the aforementioned serving systems.
9 Conclusion
@@
Speculative decoding aims to reduce inference latency by increasing computational overhead, yet this trade-off depends on the variations in request patterns, system configurations, and SLOs. We explore how these variables influence speculative decoding and have developed SpecServe, an adaptive speculative decoding framework that addresses these variations through request-level speculative length control. Experimental results demonstrate that SpecServe can effectively and consistently reduce inference latency while maintaining lower computational costs.

### v2

@@ 7.3. Ablation Study
As shown in Tab. 2, we conducted an ablation study of the performance of different dynamic methods on real-world production traces, exploring each method’s speedup sources across three dimensions: average speculative length, average accepted draft tokens, and average draft token acceptance rate. Specifically, AdaSpec- utilizes an adaptive drafter, without the confidence prior verifier.
Benefits of adaptive drafter. The adaptive drafter achieves significant performance improvements by dynamically controlling the speculative length, delivering an additional speedup ratio of 12% to 55%. Compared to traditional threshold-based methods, AdaSpec- increases the average draft token acceptance rate by 8% to 15%, benefiting from its ability to promptly terminate the draft model’s execution. The adaptive drafter can reduce the average speculative length (# Draft Tokens (Avg.)) by 2.45 tokens while only decreasing the number of actually accepted draft tokens (# Accepted Draft Tokens (Avg.)) by 0.19, indicating that AdaSpec-’s efficient acceleration primarily stems from its high conversion efficiency from draft tokens to accepted tokens. Clearly, compared to traditional threshold-based methods, the adaptive drafter employs confidence-based prediction model to precisely estimate the expected number of accepted tokens per batch. This enables finer-grained control over speculative length to improve token acceptance rates, and more efficient resource scheduling to enhance acceleration ratios.
@@
Figure 13. Fluctuations in batch sizes and speculative lengths under different methods and traces.
Benefits of confidence prior verifier. Request-level speculative length control further improves the speedup ratio. AdaSpec achieves the highest average draft token acceptance rate among all dynamic methods, which is the key reason for its superior acceleration performance. Compared to AdaSpec-, AdaSpec slightly reduces the average number of accepted draft tokens (by only 0.015 on average) while also shortening the average speculative length (by 0.14 on average), ultimately increasing the average draft token acceptance rate by up to 2%. This improvement demonstrates that the confidence prior verifier exhibits high selectivity when eliminating draft tokens, with only one out of every ten draft tokens likely being accepted, which confirms the effectiveness of the draft token elimination strategy. As described above, the confidence prior verifier serves as a more fine-grained resource scheduling mechanism, further improving token acceptance rates and speedup ratios.
7.4. Behavior Analysis
@@
As illustrated in Fig. 13, we analyzed optimal vanilla speculative decoding (with a speculative length of 3), threshold-based dynamic speculative decoding, and AdaSpec across traces from three production environments. Within a 10-minute observation period, we monitored the batch size and speculative length every 5 seconds.
Effective speculative length control. In various production trace, AdaSpec exhibits a notable dynamic adjustment characteristic between its speculative length and batch size. Specifically, when the batch size increases with fluctuating request patterns, the speculative length decreases accordingly, and vice versa. This inverse regulation mechanism ensures efficient system operation under varying load conditions. Taking a scenario with significant request rate variations (trace 1) as an example, during high request rate phases, the system adapts to larger batch sizes (fluctuating between 10 and 20) by reducing the speculative length (varying between 2 and 4). Conversely, during low request rate phases, it increases the speculative length (ranging from 4 to 7) to fully utilize smaller batch sizes (varying between 0 and 10), thereby optimizing resource efficiency based on real-time load conditions.
In contrast, threshold-based dynamic speculative decoding methods exhibit entirely different characteristics, with the speculative length consistently remaining near the preset maximum value of 8, almost unaffected by batch size variations. Although this approach simplifies the control strategy through straightforward confidence scores and threshold judgments, it results in a significantly higher average speculative length compared to AdaSpec, leading to resource wastage. Meanwhile, traditional fixed-length speculative decoding methods face even greater limitations. For instance, in most scenarios, the optimal preset speculative length (SL=3) performs similarly to AdaSpec ’s speculative length (2-4) in the first half of trace 1. However, when the batch size drops sharply in the latter half of trace 1, the fixed length of 3 fails to fully exploit the potential of speculative decoding, unlike AdaSpec (with a speculative length of 4 to 7), ultimately increasing task completion latency.
Therefore, AdaSpec ’s speculative length control mechanism demonstrates significant advantages in real-world production environments. Its ability to dynamically adjust computational resource allocation not only effectively reduces resource waste but also enhances the system’s capability to handle bursty requests. By striking an optimal balance between performance and efficiency, AdaSpec provides an exemplary approach for practical cloud service systems.
8. Related Work
Speculative decoding solutions. The efficiency of speculative decoding relies on the draft model’s speed and token acceptance rate, motivating research into fast draft models that closely approximate the target model. There are two strategies: (1) Independent draft models, which involve introducing a high-efficiency independent draft model to generate draft tokens. This model could be a variant from the same series as the target model (Leviathan et al., 2023; Sun et al., 2023), a fine-tuned model (Miao et al., 2024; Kim et al., 2023; Liu et al., 2024d; Zhou et al., 2024), or based on databases from other textual resources (He et al., 2024); (2) Self-draft models, which adjust the structure of the target model to enable it to rapidly generate multiple draft tokens (Cai et al., 2024; Li et al., [n. d.]; Fu et al., 2024a; Zhang et al., 2024; Li et al., 2024). For instance, the Multi-Token Prediction (MTP) adds multiple prediction heads to the target model, allowing it to generate consecutive draft tokens in parallel (Liu et al., 2024c). In this paradigm, the speculative length can be interpreted as the number of active prediction heads. AdaSpec offers universal adaptive control over the speculative length for both approaches, dynamically determining the optimal length for each request based on token-level confidence scores. Contemporaneous with AdaSpec, AdaServe (Li et al., 2025) also explores adaptive speculative decoding but differs: (1) AdaServe’s primary objective is to satisfy the customized SLOs of each requests, whereas AdaSpec aims to balance generation latency with SLO attainment to ensure stable acceleration. (2) AdaServe targets tree-based speculative decoding, while AdaSpec is specifically optimized for sequential speculative decoding.
Additionally, speculative decoding has been adapted for specific use cases, such as processing ultra-long or medium-length sequences, by designing specialized draft models (Sun et al., 2024a; Chen et al., 2024). This diversification of draft models has advanced the field, and AdaSpec provides fine-grained, adaptive control for any of them.
Resource-efficient LLM serving. Modern LLM serving systems use various orthogonal techniques to reduce latency and improve throughput. For instance, Sarathi-Serve (Agrawal et al., 2024) maximizes throughput with chunked-prefills and stall-free scheduling. Prefill-Decoding (PD) Disaggregation (Patel et al., 2024; Qin et al., 2025; Zhong et al., 2024) assigns prefill and decoding phases to separate GPUs to meet their respective SLOs. For multi-tenant scenarios, Llumnix (Sun et al., 2024b) optimizes cross-node resources through runtime rescheduling. AdaSpec further enhances efficiency by introducing fine-grained speculative length control, enabling speculative decoding as a stable and high-performance decoding solution. These techniques operate orthogonally, collectively optimizing resource utilization in LLM serving systems.
9. Conclusion
@@
Speculative decoding aims to reduce inference latency by increasing computational overhead, yet this trade-off depends on the variations in request patterns, system configurations, and SLOs. We explore how these variables influence speculative decoding and have developed AdaSpec, an adaptive speculative decoding framework that addresses these variations through request-level speculative length control. Experimental results demonstrate that AdaSpec can effectively and consistently reduce inference latency while maintaining lower computational costs.
Acknowledgements.
We thank the anonymous reviewers for their insightful comments that helped improve this work. We also thank Zhaoyu Fan, Bowen Han, Zehua He, and Chengyue Wu for their helpful comments. This work was supported in part by the Science and Technology Commission of Shanghai Municipality (24DP1500704 and 24YL1901100), the Fundamental Research Funds for the Central Universities (22120230311), and the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001).